Home >Documents >M . Pojer, on behalf of the usual bunch of suspects…

M . Pojer, on behalf of the usual bunch of suspects…

Date post:01-Jan-2016
Category:
View:43 times
Download:1 times
Share this document with a friend
Description:
M . Pojer, on behalf of the usual bunch of suspects…. Superconducting circuits re-commissioning after the 2011 Christmas break. Cryogenics master plan. N & w-e. 7/24. Fundamental dates. Starting from w.6 (Feb.6) the machine shall be patrolled - PowerPoint PPT Presentation
Transcript:

Slide 1

M. Pojer, on behalf of the usual bunch of suspectsSuperconducting circuits re-commissioning after the 2011 Christmas break1/24/2012#Re-commissioning after the 2011 Christmas breakCryogenics master plan

N & w-e7/241/24/2012#Re-commissioning after the 2011 Christmas breakFundamental datesStarting from w.6 (Feb.6) the machine shall be patrolledFrom Feb.6, start of ElQA on sector 67 (the day after on S23)From Feb.8, start of powering tests (Phase II directly), after cable connection and PC unlocking

Shifts proposal (experts in CCC [EPC-QPS-MP3]/OP)A shift (17:00-23:00) for the first two weeks (till Feb.19), w-e excluded / A2 shifts (M+A 7:00-15:00-23:00) for the first 2 w-e / M+A2 shifts starting from w.8 (Feb.23) / 7/24PIC on call

AM+AAM+AA+N7/24Chamonix1/24/2012#Re-commissioning after the 2011 Christmas breakPowering parameters

1/24/2012#Re-commissioning after the 2011 Christmas breakStrategy for ElQA

1/24/2012#Re-commissioning after the 2011 Christmas breakStrategy for ElQA

No specific ElQA test will be done in the tunnel as the boards will be singularly tested and transported in a safe envelope. Coherency check done by CRYO during cool-down!1/24/2012#Re-commissioning after the 2011 Christmas breakThe baseline (t.b.c. in Chamonix) is 4 TeVReconfiguration of PCs (HW and SW limits)Main circuits (6770 A required for RB; commissioned up to 6850 A)No extra powering tests / no pyramidsNo need for reconfiguration of energy extraction needed for 5 TeVInductance coefficients recalculation (1 ramp at least)No quench issue till ~5 TeV (RB.A78 had a training quench below!)RB.A78 - ElQA tested up to 1.6 kV instead of 1.9 kV due to bad insulation on magnet B30.R7 (NC 1060444) --> ok for 4 TeV, veto to be lifted for higher energies

2011 baseline

1/24/2012#Re-commissioning after the 2011 Christmas breakThe baseline (t.b.c. in Chamonix) is 4 TeVIPQ/DsNo extra powering tests (change of I_nom)No issue to go to 4 TeV (weakest magnet is RD2.R8, which is limited to 6.8 TeV)ITNormally commissioned to 5 TeVIT.R1 - weak electrical insulation of QH YT1121 to coil and/or ground (NC 1017174) --> performance of a new high capacitance PS under evaluation; verification if the energy deposited by the new DQHDS is also enough for 4 TeV ongoing (K.Dahlerup-Petersen)CorrectorsAll yearly commissioned to at least 5 TeVSome limitations on a few 120 A correctors (not for 4 TeV)Cooling problems on the 120 A correctors of ITs (agreement with ABP to commission them for use in MDs, but not needed even below 1 m b*)

EDMS doc. No11139651/24/2012#Re-commissioning after the 2011 Christmas breakProposed parameters

1/24/2012#Re-commissioning after the 2011 Christmas breakProposed parameters

1/24/2012#Re-commissioning after the 2011 Christmas breakMain changes

1/24/2012#Re-commissioning after the 2011 Christmas break

Software and toolsKajetan and Michael are working on a new GUI for follow-up, but also execution of tests. Debugging ongoing

1/24/2012#Re-commissioning after the 2011 Christmas breakExotic stuff: QPS actionsRQTD/F threshold increase doneRQ9/10 threshold has been changed but a controller upgrade still has to be doneR2E related upgrade:Automatic power cycle option in 8 units in the machine not yet doneSplice detection firmware upgrade (in units located in high radiation areas) is still an optionNew IPQ cards for radiation upgrade to be done once we have the cards (maybe on the first technical stop) no specific test more than the normal campaign is needed to validate the changeCrate controller upgrade in cell 8/9 still to be done (Reiner commented that its anyway possible to run without it)A procedure on the heaters firing at 200 V is being prepared (Knud). Markus underlined that the java program to fire the heaters will have to be tested as soon as possible. No IPQs IPDs and ITs heater will be fired during IST (its done during the commissioning step of PIC2)During the PIC2 tests, for most circuits the QPS trip signal is given by the sequencer, for some circuits this is being done by a QPS expert. Can this be done by an operator or MP3 member?F3/F4 circuit breaker: half of the machine done, still to be decided whether the other half will also be done (Reiner and Knud think we better do it)Installation of new signal cabling for the magnet and busbar voltage to the DQQDI detectors on all RQ8/9/10 circuits: a meeting will be held this afternoon to discuss itUndulator test will be organized on the spare in SM18. Special tests of the undulator in LHC: not clear, check with A.SiemkoEE system of 4-5 600 A circuits has been changed (PNO.b1 has to be done) list will be provided by KnudThe campaign to change the switches of the QHPS will be finished before the HW commissioning campaign, then Knud stressed that heaters discharge procedure is important1/24/2012#Re-commissioning after the 2011 Christmas breakPossible impact of RF re-commissioning (from P. Baudrenghien)Planning:The RF should have a first sector cold on Feb 20, and the second on Feb 22-23P. Maesen would start conditioning around Feb 22By Feb 27th we would have a few cavities conditioned and we start the setting-up of the RF loops. This will last till March 7th

Modifications during shutdown:Replacement of 8 drivers (out of 16)Replacement of 4 klystrons (out of 16)Replacement of ALL SMC cables in the UX45 faraday cage

Consequences:We will find problems when we start setting-up the RF between Feb 27th and March 7th. That will call for access in UX45 (not tunnel). Access is possible only if the magnets are at injection energy. So we will need to share time between powering tests and RF access

1/24/2012#Re-commissioning after the 2011 Christmas break

Popular Tags:
of 14/14
M. POJER, ON BEHALF OF THE USUAL BUNCH OF SUSPECTS… SUPERCONDUCTING CIRCUITS RE- COMMISSIONING AFTER THE 2011 CHRISTMAS BREAK
Embed Size (px)
Recommended